OpenStudy (anonymous):

simplify completely \[2\sqrt{5}(3\sqrt{3}+4\sqrt{2})\]

OpenStudy (espex):

\[2\sqrt{5}(3\sqrt{3}+4\sqrt{2}) \rightarrow (2 \times 3 \times \sqrt{5}\times \sqrt{3}) + (2 \times 4 \times \sqrt{5} \times \sqrt{2}) \rightarrow 6\sqrt{5 \times 3} + 8\sqrt{5 \times 2}\] \[6\sqrt{5 \times 3} + 8\sqrt{5 \times 2} \rightarrow 6\sqrt{15}+8\sqrt{10}\]